What will be the new area of 400cm square of a solid if it's temperature is raised by 10kelvin
1
Expert's answer
2020-06-17T09:25:52-0400

If the linear expansion coefficient is "\alpha" , the new area after heating will be


"A=A_0(1+\alpha T)^2=400(1+10\alpha)^2\text{ cm}^2."

For instance, for brass square with "\alpha=19\cdot10^{-6}\text{ K}^{-1}" the new area will be 400.15 cm2